A company is considering constructing a plant to manufacture a proposed new product. The land costs $300,000, the building costs $600,000, the equipment costs $250,000, and $100,000 additional working capital is required. It is expected that the product will result in sales of $750,000 per year for 10 years,
1
Expert's answer
2021-10-04T02:30:59-0400

Modified B-C ratio = (PW of annual benefits - PW of annual expenses) / (Land + Building + Equipment + Working capital - PW of salvage value)

"= [(750,000 - 475,000) x P/A(15%, 10)] / [300,000 + 600,000 + 250,000 + 100,000 - (400,000 + 350,000 + 50,000) x P/F(15%, 10)]\\

= (275,000 x 5.0188**) / [1,250,000 - (800,000 x 0.2472**)]\\

= 1,380,170 / (1,250,000 - 197,760)\\

= 1,380,170 / 1,052,240\\

= 1.31\\"

Since Modified B-C Ratio is greater than 1, Project is acceptable